# Federated Continual Instruction Tuning

### v1

@@ Abstract
A vast amount of instruction tuning data is crucial for the impressive performance of Large Multimodal Models (LMMs), but the associated computational costs and data collection demands during supervised fine-tuning make it impractical for most researchers. Federated learning (FL) has the potential to leverage all distributed data and training resources to reduce the overhead of joint training. However, most existing methods assume a fixed number of tasks, while in real-world scenarios, clients continuously encounter new knowledge and often struggle to retain old tasks due to memory constraints. In this work, we introduce the Federated Continual Instruction Tuning (FCIT) benchmark to model this real-world challenge. Our benchmark includes two realistic scenarios, encompassing four different settings and twelve carefully curated instruction tuning datasets. To address the challenges posed by FCIT, we propose dynamic knowledge organization to effectively integrate updates from different tasks during training and subspace selective activation to allocate task-specific output during inference. Extensive experimental results demonstrate that our proposed method significantly enhances model performance across varying levels of data heterogeneity and catastrophic forgetting. Our source code and dataset will be made publicly available.
@@ 1 Introduction
Large Multimodal Models (LMMs) [2, 29, 30], which integrate Large Language Model [4, 18, 44] with a visual encoder and multimodal projector to bridge visual and textual modalities, have exhibited impressive visual understanding and complex reasoning abilities. A crucial factor in this success is the supervised fine-tuning of LMMs using huge and diverse visual instruction-following data [29] to align with human preferences. However, collecting such vast amounts of training data and computational resources for joint fine-tuning is impractical for most researchers. Federated Learning (FL) [34, 26, 39], as a decentralized paradigm, offers a viable alternative by leveraging distributed data and computational resources for local training while integrating local weights to produce a unified model. This paradigm accommodates constraints on storage and computation while ensuring privacy protection.
Generally, most existing FL frameworks [50, 55, 7] are modeled in static, closed-world scenarios [56, 3], where a fixed set of tasks is predefined and unchanged. However, real-world applications are dynamic [59, 53], requiring models to continuously acquire new knowledge while retaining previously learned tasks. Taking a realistic health emergency as an example, different hospitals act as local clients when a major disease outbreak occurs. Large hospitals can leverage FL to collaboratively train on their case data, building a comprehensive virus knowledge base, while smaller clinics update their own cases using this global knowledge. Over time, both large and small hospitals need to keep updating the global knowledge base through continual learning methods to cope with the event. Additionally, the emergence of new strains necessitates simultaneous integration of knowledge and response strategies to minimize potential losses. In this scenario, traditional FL methods struggle with newly arrived knowledge, while Continual Learning (CL) methods alone do not facilitate knowledge sharing between clients. Only an organic integration of both can address this real-world problem. Recently, numerous Federated Continual Learning methods have emerged to addressed this challenge in traditional image classification tasks [11, 12, 54, 51]. However, their methodological and task-setting limitations make them insufficient for current LMMs applications. Therefore, a more comprehensive benchmark is needed to better simulate the practical application of LMMs in real-world scenarios.
@@
In this work, we first establish a Federated Continual Instruction Tuning (FCIT) benchmark to fill this gap. Specifically, we simulate two client-level realistic scenarios: (i) Homogeneous FCIT refers to an FL system learning a series of instruction tuning data sequentially, with different clients learning the same task at each stage. (ii) Heterogeneous FCIT, on the other hand, requires the FL system to collaborate on different tasks simultaneously, with different clients potentially learning different tasks in the same stage. Building on this, we define two settings of dataset for each scenario: Capability-related and Task-related. The former evaluates the model’s ability to integrate multiple dimensions of instruction following in a short period, while the latter assesses the model’s performance during long phases of continual learning. For dataset selection, we curate twelve instruction tuning datasets unseen during LMM fine-tuning or with low zero-shot performance, preventing information leakage [21, 13]. In addition, we introduce varying degrees of data heterogeneity for each setting to challenge the model’s performance in the non-IID situation [25]. An illustration of our FCIT benchmark is provided in Figure 1. To the best of our knowledge, this is the first work to introduce a comprehensive benchmark for federated learning of LMM in a continual learning setting.
@@
To effectively address the challenges posed by FCIT, we propose a noval Dynamic knowledge organIzation and Subspace seleCtive activatiOn (DISCO) framework. Specifically, we identify FCIT challenges into two types: conflicts between different tasks within the same stage and conflicts between old and new tasks across different stages. The former can cause catastrophic forgetting by altering the parameter space of previous tasks when learning new ones without access to past data, while the latter requires the model to integrate and organize knowledge from different tasks to derive unified representations. Therefore, we first propose Dynamic Knowledge Organization (DKO), which leverages a dynamic cache at the global server to store task-specific parameters. Using a unique identity token matching mechanism, it systematically organizes knowledge for different tasks into corresponding subspaces within the cache, effectively mitigating two types of conflicts. To better utilize the organized task subspaces in the dynamic cache, we introduce Subspace Selective Activation (SSA), which selectively activates subspaces relevant to the test input while filtering out irrelevant outputs, leading to significant performance improvements. Consequently, these designs enable our framework to efficiently tackle the data heterogeneity and catastrophic forgetting in FCIT. In summary, our major contributions are:
@@
  • We propose a noval DISCO framework that integrates dynamic knowledge organization and subspace selective activation to efficiently address data heterogeneity and catastrophic forgetting in FCIT settings.
@@ 2 Related Work
Federated Continual Learning. In traditional vision field, Federated Continual Learning (FCL) [49] aims to adapt the global model to new data while maintaining the knowledge of the old task. From the perspective of LMMs, researches in this setting face two main challenges: (1) Most of the methods are designed for traditional vision tasks (i.e., image classification). For instance, MFCL [1] employs a generative model to synthesize images of previously learned classes, thereby mitigating forgetting. PILoRA [12] introduces a prototype re-weight module to address the classifier bias caused by data heterogeneity and obtain unified knowledge through LoRA [16] fusion. Despite progress in classification tasks, complex designs remain challenging to adapt for LMMs research. (2) FCL follows a single data composition by splitting a dataset (e.g. ImageNet [10]) into different tasks based on classes, whereas LMMs face greater challenges in continual learning due to their diverse tasks with varying styles [6, 52, 5].
@@ 3.1 Preliminaries
where $\theta$ denotes the trainable parameters during fine-tuning, $\textbf{x}_{res,<i}$ is the response tokens from all previous turns before the current prediction token $x_{i}$. Then, the loss function of fine-tuning LMMs can be expressed as:
$\mathcal{L}_{\theta}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i=1}^{L}\log p_{\theta}(x_{i}^{j}|\textbf{x}_{v}^{j},\textbf{x}_{ins}^{j},\textbf{x}_{res,<i}^{j}).$ (2)
@@
Continual learning aims to minimize the loss on the current task while retaining knowledge from previous tasks. Formally, given a sequence of datasets $\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{T}$, the optimization objective of at task $t$ is:
@@ 3.2 Federated Continual Instruction Tuning
  • Task-related. To evaluate the performance of different methods in a long-phase continual learning situations, we selected 8 datasets: ImageNet-R, ArxivQA, IconQA, CLEVR-Math, OCR-VQA, Flickr30k, FigureQA, and super-CLEVR as distinct stages of continual learning.
@@ 4.1 Dynamic Knowledge Organization
$\Delta\textbf{W}=\textbf{B}\textbf{A}\Leftrightarrow\underbrace{\textbf{B}_{1}\textbf{A}_{1}+\cdots+\textbf{B}_{T}\textbf{A}_{T}}_{\text{task-specific subspace of each task}},$ (5)
@@
Inspired by the widespread use of prototypes in FL fields [42, 17, 43], We propose to distinguish knowledge across tasks by using the feature mean of each client training data as an Identity Token. Specifically, considering the uniqueness of textual inputs in the visual instruction tuning task, we introduce a text encoder $f_{ins}$ for each client to extract the feature of the input instruction $\textbf{x}_{ins,j}^{t}$ during training, and take the mean value $\mathbf{\mu}_{k}^{t}$ as the local identity token of the $k$-th client on the task $t$ at the end of training:
@@
where $n_{k}^{t}$ is the number of training samples of client $k$ at task $t$ and we use CLIP’s text encoder [37] in our experiments. Notably, the text encoder $f_{ins}$ remains frozen during training, which prevents additional training overhead.
Using the identify tokens, we first apply cosine similarity along with a threshold to match the uploaded local identify token $\mathbf{\mu}_{j}^{t}$ with the global identify token $\mathbf{\tilde{z}}_{i}$:
@@
where $\tau$ denotes a pre-defined threshold. For paired local identity tokens, we update the corresponding global identify token. For mismatched local identity tokens, we pair them two by two using Eq. (7) and initialize new global identity tokens. This process can be formalized as:
@@
where $m$, $n_{\mathbf{\tilde{z}}_{i}}$ and $n_{j}^{t}$ denote the number of matched local identify tokens, the number of samples used to form the previous $i$-th global identity token, and the number of samples from the local client $j$ at task $t$, respectively. Then, we leverage this matching process to guide the update or initialize of task-specific subspace $\{\mathbf{\theta}_{i}=\textbf{B}_{i}\textbf{A}_{i}\}$ in Eq. (5):
@@ 4.2 Subspace Selective Activation
where $\mathcal{W}\in\mathbb{R}^{r\times r}$ is typically a identity matrix and can thus be omitted. In this paper, we redefine $\mathcal{W}$ as the activation matrix that dynamically responds to the test input. Specifically, we treat $\mathcal{W}$ as the product of an identity matrix and an activation factor (i.e., $\mathcal{W}=\alpha\cdot\mathbf{I}_{r\times r}$), where $\alpha=1$ denotes full activation, and $\alpha=0$ means that the output is fully masked. Therefore, for each subspace $\{\textbf{B}_{1}\textbf{A}_{1},\cdots,\textbf{B}_{T}\textbf{A}_{T}\}$ in dynamic cache, we can control the activation or inhibition of the corresponding output by adjusting its activation factor $\{\alpha_{1},\cdots,\alpha_{T}\}$.
where $\varepsilon$ is the temperature coefficient.
Overall, our proposed SSA can be formulated as follows:
$\Delta\textbf{W}$ $=\bar{\textbf{B}}\bar{\mathcal{W}}\bar{\textbf{A}}$ (13)
5 Experiments
5.1 Experimental Setup
Datasets. The dataset composition is detailed in Section 3.2. For the experimental setup, we implement two types of dataset-level setting: Capability-related and Task-related, based on the two client-level realistic scenarios, Hom-FCIT and Het-FCIT, respectively. At each stage, the own data of local clients are partitioned according to the Dirichlet distribution [25], with three partition $\beta$ for each setting to model different levels of data heterogeneity. More details are provided in Appendix A.
@@ 5.1 Experimental Setup
Baselines & Evaluation Metrics. We compare our meth-od with continual learning methods such as LwF [27], EWC [22], L2P [46], O-LoRA [45], MoELoRA [6], and also with the LoRA merging method from the federated continual learning approach PILoRA [12], referred to as M-LoRA. We follow a rehearsal-free continual learning setting [58], where only the data for the current task is available. All comparison methods are carefully calibrated to ensure fairness of evaluations.
@@ 5.2 Main Results
Results are shown in Table 1Table 2Figure 4, and Figure 5. Under Hom-FCIT setting, our proposed DISCO achieves the best performance in both capability-related and task-related dataset setting, surpassing the best baseline by an average of 4.84% and 1.43% in the Last and Avg metrics, respectively. As illustrated in Figure 4, DISCO effectively mitigates the forgetting of previous tasks while learning new ones, outperforming other methods in the challenging long-stage continual learning setting.
@@ 5.3 Ablation Study
The detailed ablation studies are provided in Table 3. In this paper, we use CLIP’s text encoder to extract textual feature as identity tokens. This section tests two alternatives: using only visual features from CLIP’s visual encoder, and combining visual and textual features. From Table 3a, we observe that both alternatives degrade performance. This is potential because, in visual instruction tuning datasets, task similarities at the image level are higher than at the textual level (e.g. CLEVR-Math and super-CLEVR), making visual features less effective than textual ones.
@@
We also test the location of LoRA embedding. As can be seen in Table 3c, embedding in the attention layer alone proved less effective than in the FFN layer. Moreover, embedding LoRA in every linear layer did not offer significant improvement and resulted in higher parameter transmission. Therefore, we chose to embed LoRA only in the FFN layer as a balanced solution. We provide more ablation and visualization results in Appendix B

### v2

@@ Abstract
A vast amount of instruction tuning data is crucial for the impressive performance of Large Multimodal Models (LMMs), but the associated computational costs and data collection demands during supervised fine-tuning make it impractical for most researchers. Federated learning (FL) has the potential to leverage all distributed data and training resources to reduce the overhead of joint training. However, most existing methods assume a fixed number of tasks, while in real-world scenarios, clients continuously encounter new knowledge and often struggle to retain old tasks due to memory constraints. In this work, we introduce the Federated Continual Instruction Tuning (FCIT) benchmark to model this real-world challenge. Our benchmark includes two realistic scenarios, encompassing four different settings and twelve carefully curated instruction tuning datasets. To address the challenges posed by FCIT, we propose a dynamic knowledge organization to effectively integrate updates from different tasks during training and subspace selective activation to allocate task-specific output during inference. Extensive experimental results demonstrate that our proposed method significantly enhances model performance across varying levels of data heterogeneity and catastrophic forgetting. Code and dataset are released at https://github.com/Ghy0501/FCIT.
@@ 1 Introduction
Large Multimodal Models (LMMs) [2, 32, 34], which integrate Large Language Model [4, 20, 49, 46] with a visual encoder and multimodal projector to bridge visual and textual modalities, have exhibited impressive visual understanding and complex reasoning abilities. A crucial factor in this success is the supervised fine-tuning of LMMs using huge and diverse visual instruction-following data [32] to align with human preferences. However, collecting such vast amounts of training data and computational resources for joint fine-tuning is impractical for most researchers. Federated Learning (FL) [38, 28, 43, 31], as a decentralized paradigm, offers a viable alternative by leveraging distributed data and computational resources for local training while integrating local weights to produce a unified model. This paradigm accommodates constraints on storage and computation while ensuring privacy protection.
@@
Generally, most existing FL frameworks [55, 58, 7] are modeled in static, closed-world scenarios [59, 3], where a fixed set of tasks is predefined and unchanged. However, real-world applications are dynamic [63, 59], requiring models to continuously acquire new knowledge while retaining previously learned tasks. Taking a realistic health emergency as an example, different hospitals act as local clients when a major disease outbreak occurs. Large hospitals can leverage FL to collaboratively train on their case data, building a comprehensive virus knowledge base, while smaller clinics update their own cases using this global knowledge. Over time, both large and small hospitals need to keep updating the global knowledge base through continual learning methods to cope with the event. Additionally, the emergence of new strains necessitates the simultaneous integration of knowledge and response strategies to minimize potential losses. In this scenario, traditional FL methods struggle with newly arrived knowledge, while Continual Learning (CL) [15, 14, 56, 60, 33] methods alone do not facilitate knowledge sharing between clients. Only an organic integration of both can address this real-world problem. Recently, numerous Federated Continual Learning methods have emerged to address this challenge in traditional image classification tasks [11, 12, 57]. However, their methodological and task-setting limitations make them insufficient for current LMM applications. Therefore, a more comprehensive benchmark is needed to better simulate the practical application of LMMs in real-world scenarios.
@@
In this work, we first establish a Federated Continual Instruction Tuning (FCIT) benchmark to fill this gap. Specifically, we simulate two client-level realistic scenarios: (i) Homogeneous FCIT refers to an FL system learning a series of instruction tuning data sequentially, with different clients learning the same task at each stage. (ii) Heterogeneous FCIT, on the other hand, requires the FL system to collaborate on different tasks simultaneously, with different clients potentially learning different tasks in the same stage. Building on this, we define two settings of datasets for each scenario: Capability-related and Task-related. The former evaluates the model’s ability to integrate multiple dimensions of instruction following in a short period, while the latter assesses the model’s performance during long phases of continual learning. For dataset selection, we curate twelve instruction tuning datasets unseen during LMM fine-tuning or with low zero-shot performance, preventing information leakage [23, 13]. In addition, we introduce varying degrees of data heterogeneity for each setting to challenge the model’s performance in the non-IID situation [27]. An illustration of our FCIT benchmark is provided in Figure 1. To the best of our knowledge, this is the first work to introduce a comprehensive benchmark for federated learning of LMM in a continual learning setting.
@@
To effectively address the challenges posed by FCIT, we propose a novel Dynamic knowledge organIzation and Subspace seleCtive activatiOn (DISCO) framework. Specifically, we identify FCIT challenges into two types: conflicts between different tasks within the same stage and conflicts between old and new tasks across different stages. The former can cause catastrophic forgetting by altering the parameter space of previous tasks when learning new ones without access to past data, while the latter requires the model to integrate and organize knowledge from different tasks to derive unified representations. Therefore, we first propose Dynamic Knowledge Organization (DKO), which leverages a dynamic cache at the global server to store task-specific parameters. Using a unique identity token matching mechanism, it systematically organizes knowledge for different tasks into corresponding subspaces within the cache, effectively mitigating two types of conflicts. To better utilize the organized task subspaces in the dynamic cache, we introduce Subspace Selective Activation (SSA), which selectively activates subspaces relevant to the test input while filtering out irrelevant outputs, leading to significant performance improvements. Consequently, these designs enable our framework to efficiently tackle the data heterogeneity and catastrophic forgetting in FCIT. In summary, our major contributions are:
@@
  • We propose a novel DISCO framework that integrates dynamic knowledge organization and subspace selective activation to efficiently address data heterogeneity and catastrophic forgetting in FCIT settings.
@@ 2 Related Work
Federated Continual Learning. In classical vision tasks, Federated Continual Learning (FCL) [54] aims to adapt the global model to new data while maintaining the knowledge of the old task. From the perspective of LMMs, research in this setting faces two main challenges: (1) Most of the methods are designed for traditional vision tasks (i.e., image classification). For instance, MFCL [1] employs a generative model to synthesize images of previously learned classes, thereby mitigating forgetting. PILoRA [12] introduces a prototype re-weight module to address the classifier bias caused by data heterogeneity and obtain unified knowledge through LoRA [18] fusion. Despite progress in classification tasks, complex designs remain challenging to adapt for LMMs research. (2) FCL follows a single data composition by splitting a dataset (e.g. ImageNet [10]) into different tasks based on classes, whereas LMMs face greater challenges in continual learning due to their diverse tasks with varying styles [6, 56, 5].
@@ 3.1 Preliminaries
where $\theta$ denotes the trainable parameters during fine-tuning, $\textbf{x}_{\text{res},<i}$ denotes all response tokens preceding the current prediction token $x_{i}$. Then, the loss function of fine-tuning LMMs can be expressed as:
$\mathcal{L}_{\theta}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i=1}^{L_{j}}\log p_{\theta}(x_{i}^{j}|\textbf{x}_{v}^{j},\textbf{x}_{ins}^{j},\textbf{x}_{res,<i}^{j}).$ (2)
@@
Continual learning aims to minimize the loss on the current task while retaining knowledge from previous tasks. Formally, given a sequence of datasets $\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{T}$, the optimization objective at task $t$ is:
@@ 3.2 Federated Continual Instruction Tuning
  • Task-related. To evaluate the performance of different methods in a long-phase continual learning situation, we selected 8 datasets: ImageNet-R, ArxivQA, IconQA, CLEVR-Math, OCR-VQA, Flickr30k, FigureQA, and super-CLEVR as distinct stages of continual learning.
@@ 4.1 Dynamic Knowledge Organization
$\Delta\textbf{W}=\textbf{B}\textbf{A}\Leftrightarrow\underbrace{\{\textbf{B}_{1}\textbf{A}_{1},\cdots,\textbf{B}_{T}\textbf{A}_{T}\}}_{\text{task-specific subspace}},$ (5)
@@
Inspired by the widespread use of prototypes in FL fields [47, 19, 48], we propose to distinguish knowledge across tasks by using the feature mean of each client’s training data as an Identity Token. Specifically, considering the uniqueness of textual inputs in the visual instruction tuning task, we introduce a text encoder $f_{ins}$ for each client to extract the feature of the input instruction $\textbf{x}_{ins,j}^{t}$ during training, and take the mean value $\mathbf{\mu}_{k}^{t}$ as the local identity token of the $k$-th client on the task $t$ at the end of training:
@@
where $n_{k}^{t}$ is the number of training samples of client $k$ at task $t$, and we use CLIP’s text encoder [41] in our experiments. Notably, the text encoder $f_{ins}$ remains frozen during training, which prevents additional training overhead.
Using the identity tokens, we first apply cosine similarity along with a threshold to match the uploaded local identity token $\mathbf{\mu}_{j}^{t}$ with the global identity token $\mathbf{\tilde{z}}_{i}$:
@@
where $\tau$ denotes a pre-defined threshold. For paired local identity tokens, we update the corresponding global identity token. For mismatched local identity tokens, we pair them two by two using Eq. (7) and initialize new global identity tokens. This process can be formalized as:
@@
where $m$, $n_{\mathbf{\tilde{z}}_{i}}$ and $n_{j}^{t}$ denote the number of matched local identity tokens, the number of samples used to form the previous $i$-th global identity token, and the number of samples from the local client $j$ at task $t$, respectively. Then, we leverage this matching process to guide the update or initialization of task-specific subspace $\{\mathbf{\theta}_{i}=\textbf{B}_{i}\textbf{A}_{i}\}$ in Eq. (5):
@@ 4.2 Subspace Selective Activation
where $\mathcal{W}\in\mathbb{R}^{r\times r}$ is typically an identity matrix and can thus be omitted. In this paper, we redefine $\mathcal{W}$ as the activation matrix that dynamically responds to the test input. Specifically, we treat $\mathcal{W}$ as the product of an identity matrix and an activation factor (i.e., $\mathcal{W}=\alpha\cdot\mathbf{I}_{r\times r}$), where $\alpha=1$ denotes full activation, and $\alpha=0$ means that the output is fully masked. Therefore, for each subspace $\{\textbf{B}_{1}\textbf{A}_{1},\cdots,\textbf{B}_{T}\textbf{A}_{T}\}$ in dynamic cache, we can control the activation or inhibition of the corresponding output by adjusting its activation factor $\{\alpha_{1},\cdots,\alpha_{T}\}$.
@@
where $\varepsilon$ is the temperature coefficient.
Overall, our proposed SSA can be formulated as follows:
$\Delta\textbf{W}$ $=\bar{\textbf{B}}\bar{\mathcal{W}}\bar{\textbf{A}}$ (13)
5 Experiments
5.1 Experimental Setup
Datasets. The dataset composition is detailed in Section 3.2. For the experimental setup, we implement two types of dataset-level settings: Capability-related and Task-related, based on the two client-level realistic scenarios, Hom-FCIT and Het-FCIT, respectively. At each stage, the own data of local clients is partitioned according to the Dirichlet distribution [27], with three partitions $\beta$ for each setting to model different levels of data heterogeneity. More details are provided in Appendix A.
Baselines & Evaluation Metrics. We compare our method with continual learning methods such as LwF [29], EWC [24], L2P [51], O-LoRA [50], MoELoRA [6], and also with the LoRA merging method from the federated continual learning approach PILoRA [12], referred to as M-LoRA. We follow a rehearsal-free continual learning setting [62], where only the data for the current task is available. All comparison methods are carefully calibrated to ensure the fairness of evaluations.
@@ 5.2 Main Results
Results are shown in Table 1Table 2Figure 4, and Figure 5. Under Hom-FCIT setting, our proposed DISCO achieves the best performance in both capability-related and task-related dataset settings, surpassing the best baseline by an average of 4.84% and 1.43% in the Last and Avg metrics, respectively. As illustrated in Figure 4, DISCO effectively mitigates the forgetting of previous tasks while learning new ones, outperforming other methods in the challenging long-stage continual learning setting.
@@ 5.3 Ablation Study
The detailed ablation studies are provided in Table 3. In this paper, we use CLIP’s text encoder to extract textual features as identity tokens. This section tests two alternatives: using only visual features from CLIP’s visual encoder, and combining visual and textual features. From Table 3a, we observe that both alternatives degrade performance. This is potential because, in visual instruction tuning datasets, task similarities at the image level are higher than at the textual level (e.g. CLEVR-Math and super-CLEVR), making visual features less effective than textual ones.
@@
We also test the location of LoRA embedding. As can be seen in Table 3c, embedding in the attention layer alone proved less effective than in the FFN layer. Moreover, embedding LoRA in every linear layer did not offer significant improvement and resulted in higher parameter transmission. Therefore, we choose to embed LoRA only in the FFN layer as a balanced solution. We provide more ablation and visualization results in Appendix B.